OpenStudy (anonymous):

The lengths of two sides of a triangle are shown below: Side 1: 3x^2 - 2x - 1 Side 2: 9x + 2x^2 - 3 The perimeter of the triangle is 5x^3 + 4x^2 - x - 3. Part B: What is the length of the third side of the triangle?

OpenStudy (anonymous):

Anyone? :/

OpenStudy (aum):

Perimeter = length of side 1 + length of side 2 + length of side 3 length of side 3 = Perimeter - length of side 1 - length of side 2

OpenStudy (anonymous):

Okay

OpenStudy (aum):

length of side 3 = Perimeter - length of side 1 - length of side 2 = Perimeter - ( length of side 1 + length of side 2 ) Add the first two expressions given in the problem and then subtract it from the perimeter.

OpenStudy (triciaal):



OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

is it an equilateral ,isosceles, or scalene triangle

OpenStudy (triciaal):

Perimeter = sum of all 3 sides 5x^3 + 4x^2 - x - 3 = 3x^2 - 2x - 1 + 9x + 2x^2 - 3 + y solve for y the length of the 3rd side

OpenStudy (triciaal):

equilateral = all 3 sides are equal isosceles = 2 sides are equal scalene = no sides are equal; all different lengths

OpenStudy (anonymous):

They don't clarify what it is, it's just a triangle..

OpenStudy (anonymous):

So, we have to subtract the perimeter by side 1 and side 2?

OpenStudy (anonymous):

@triciaal ??

OpenStudy (anonymous):

Someone?

OpenStudy (triciaal):

you can tell from the lengths subtract side 1 and side 2 from the perimeter to find the length of the 3rd side

OpenStudy (anonymous):

Okay, can you take me through the subtraction?

OpenStudy (triciaal):

5x^3 + 4x^2 - x - 3 = (3x^2 - 2x - 1 + 9x + 2x^2 - 3) + y 5x^3 + 4x^2 - x - 3 -(3x^2 - 2x - 1 + 9x + 2x^2 - 3) = y

OpenStudy (triciaal):

5x^3 + 4 x^2 - x - 3 - 3 x^2 + 2 x + 1 - 9 x - 2 x^2 + 3 = y

OpenStudy (triciaal):

now group like terms 5 x^3 + x^2 ( 4 -3 -2) +x(-1 + 2 -9) +(-3 + 1 + 3 ) = y 5 x^3 - x^2 -8 x + 1

OpenStudy (anonymous):

Okay, then thats it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Oh! So the third side is 5 x^3 - x^2 -8 x + 1 ?

OpenStudy (triciaal):

yes that's what I got

OpenStudy (anonymous):

Okay thanks a ton!!!!!!!

OpenStudy (triciaal):

if 2 + 3 + x = 7 how would you find x? same principle x = 7 - (2 + 3)

OpenStudy (anonymous):

@triciaal your great at this. I got the same thing

OpenStudy (anonymous):

One last thing, does this show that polynomials are closed under addition and subtraction?

OpenStudy (triciaal):

@lacrosseplayer22 thanks @LearningIsAwesome you are welcome

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

@triciaal thanks for being so smart and an expert @LearningIsAwesome great jjob!

OpenStudy (anonymous):

And thats because when a plynomial is added or subtracted it equals another polynomial right?

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@triciaal is on a different problem

OpenStudy (anonymous):

Well @lacrosseplayer22 do you know?

OpenStudy (anonymous):

that depends

OpenStudy (anonymous):

if its added or subtracting with an equal value on both sides yes

OpenStudy (anonymous):

am I right @triciaal or not...or somewhat...

OpenStudy (anonymous):

Okay! thanks SO much to both of you! Especially @triciaal :)

OpenStudy (anonymous):

wait a sec. I mat not be right. im waiting for @triciaal 's response

OpenStudy (anonymous):

I think it shows closure because when a polynomial is subtracted or added it equals another polynomial, right?

OpenStudy (triciaal):

an equation is an equation. a + b = c c-b=a and c - a = b I am looking for the textbook version

OpenStudy (anonymous):

Oh, so they are closed!